OpenStudy (anonymous):

what is the square root of 75x^2y

OpenStudy (anonymous):

ty

OpenStudy (saifoo.khan):

Factor 75.

OpenStudy (anonymous):

hay how do u fined where the answer is ?

OpenStudy (saifoo.khan):

\[\sqrt{75} = \sqrt{5 \times 5 \times 3} = 5 \sqrt3\]

OpenStudy (anonymous):

ooooo ty can u help me with one more answer?

OpenStudy (saifoo.khan):

Therefore, \[5x \sqrt{3y}\]

OpenStudy (saifoo.khan):

Sure. Throw it.

OpenStudy (anonymous):

what is 147m^3n^3

OpenStudy (saifoo.khan):

Is there a radical or something?

OpenStudy (anonymous):

no they just want us to find out what the square root of the number is and put the m and nsomewhere im having some troulbe understanding it really

OpenStudy (saifoo.khan):

Can you factor 147?

OpenStudy (anonymous):

147 can be factored into 3 by 49

OpenStudy (saifoo.khan):

perfect.

OpenStudy (saifoo.khan):

\[\sqrt{147} = \sqrt{3 \times 49} = \sqrt{3 \times 7 \times 7} = 7 \sqrt3\]

OpenStudy (anonymous):

now what do i do with the m^3 and N^3

OpenStudy (saifoo.khan):

Look, Break it into exponent two. \[\Large \sqrt{m^3} = \sqrt{m^2 \times m} = m \sqrt m\]

OpenStudy (saifoo.khan):

Same as the case with n^3

OpenStudy (anonymous):

so it would be 7mn square root and 3mn?

OpenStudy (saifoo.khan):

Omg! you are a guru in this!

OpenStudy (anonymous):

ty so much by the way i gtg

OpenStudy (saifoo.khan):

Welcome. No worries. Be back soon! ^_^